two numbers with a sum of -4 and a difference of 10
Answer by checkley77(12844)

X+Y=-4
X-Y=10 ADD
-------
2X=6
X=6/2
X=3 ANS.
3+Y=-4
Y=-4-3
Y=-7 ANS.
3-(-7)=10
3+7=10
10=10